If y=x (to the power 7÷2) find dy÷dx
1
Expert's answer
2018-08-06T10:57:08-0400
y=x^(7/2 )
Apply the Power Rule:
dy/dx (x^a )=a∙x^(a-1)
So,
dy/dx (x^(7/2 ) )=7/2∙x^(7/2-1)=7/2 x^(5/2)